OpenStudy (anonymous):

DESCRIBE A PATTERN IN THE SEQUENCE OF NUMBERS. WHAT IS THE NEXT NUMBER 1,11,121,1331

jimthompson5910 (jim_thompson5910):

You're multiplying by 11 each time So the next number is 1331*11 = 14641

OpenStudy (anonymous):

you can find the nth term of the sequence using\[a_{n}=11^{n-1}\]

OpenStudy (anonymous):

for example the third term we know is 121, so\[a _{3}=11^{3-1}=11^2=121\]

OpenStudy (phi):

Or Pascal's triangle, which is the same as multiplying by 11

OpenStudy (anonymous):

cool

OpenStudy (phi):

except the triangle has 10's in it. 1 5 10 10 5 1

OpenStudy (anonymous):

@jahed, this is pascal's triangle